# The Complexity of Resilience Problems via Valued Constraint Satisfaction

### v5

@@ 8. Application: Resilience
Let $\tau$ be a finite relational signature and $\mu$ a conjunctive query over $\tau$. The input to the resilience problem for $\mu$ consists of a finite $\tau$-structure ${\mathfrak{A}}$, called a ($\tau$-)database [^1] To be precise, a finite relational structure is not exactly the same as a database, because the latter may not contain elements that are not contained in any relation. This difference, however, is inessential for the problems studied in this article., and the task is to compute the number of tuples that have to be removed from relations of ${\mathfrak{A}}$ so that ${\mathfrak{A}}$ does not satisfy $\mu$. This number is called the resilience of ${\mathfrak{A}}$ (with respect to $\mu$). As usual, this can be turned into a decision problem where the input also contains a natural number $u\in{\mathbb{N}}$ and the question is whether the resilience is at most $u$. Clearly, ${\mathfrak{A}}$ does not satisfy $\mu$ if and only if its resilience is 00.
@@
A natural variation of the problem is that the input database is a bag database, meaning that it may contain tuples with multiplicities. Formally, a multiset relation on a set $A$ of arity $k$ is a multiset with elements from $A^{k}$ and a bag database ${\mathfrak{A}}$ over a relational signature $\tau$ consists of a finite domain $A$ and for every $R\in\tau$ of arity $k$, a multiset relation $R^{\mathfrak{A}}$ of arity $k$. A bag database ${\mathfrak{A}}$ satisfies a union of conjunctive queries $\mu$ if the relational structure obtained from ${\mathfrak{A}}$ by forgetting the multiplicities of tuples in its relations satisfies $\mu$. In this article, we focus on bag databases whose relevance has already been discussed in the introduction.
@@
The basic resilience problem defined above can be generalized by admitting the decoration of databases with a subsignature $\sigma\subseteq\tau$, in this way declaring all tuples in $R^{{\mathfrak{A}}}$, $R\in\sigma$, to be exogenous. This means that we are not allowed to remove such tuples from ${\mathfrak{A}}$ to make $\mu$ false; the tuples in the other relations are then called endogenous. For brevity, we also refer to the relations in $\sigma$ as being exogenous and those in $\tau\setminus\sigma$ as being endogenous. If not specified, then $\sigma=\emptyset$, i.e., all tuples are endogenous. As an alternative, one may also declare individual tuples as being endogenous or exogenous. Under bag semantics, however, this case can be reduced to the one studied here (see Remark 8.21). The resilience problem that we study is summarized in Figure LABEL:fig:res-problem. Note that for every $\mu$ this problem is in NP. We remark that in Section 8.3, we briefly deviate from our focus on unions of conjunctive queries and, by adapting our results from Section 8.2, we establish a complexity dichotomy for resilience problems for two-way regular path queries; all the necessary definitions will be given in Section 8.3.
@@
Observe that a finite $\tau$-structure satisfies $\mu$ if and only if it does not have a homomorphism to the $\tau$-structure ${\mathfrak{B}}$ with domain $B=\{0,1\}$ and the relations $R^{{\mathfrak{B}}}=\{(0,1),(1,1)\}$ and $S^{{\mathfrak{B}}}=\{(0,0),(0,1)\}$ (see Figure 2). We turn ${\mathfrak{B}}$ into the valued structure $\Gamma$ with domain $\{0,1\}$ where $R^{\Gamma}(0,1)=R^{\Gamma}(1,1)=0=S^{\Gamma}(0,0)=S^{\Gamma}(0,1)$ and $R^{\Gamma}$ and $S^{\Gamma}$ take value 1 otherwise. Then $\operatorname{VCSP}(\Gamma)$ is precisely the resilience problem for $\mu$ (with multiplicities). Our results reprove the result from (Makhija and Gatterbauer, 2023) that even with multiplicities, the problem can be solved in polynomial time (see Theorem 7.18, Proposition 8.20, and Example 8.12).
Figure 2. The query $\mu$ from Example 8.2 (on the left) and the corresponding structure ${\mathfrak{B}}$ (on the right).
[…]
Arrow depiction of the query from the example and the dual structure.
@@
This query is linear in the sense of Freire, Gatterbauer, Immerman, and Meliou and thus its resilience problem without multiplicities can be solved in polynomial time (Theorem 4.5 in (Meliou et al., 2010); also see Fact 3.18 in (Freire et al., 2015a)). Our results reprove the result from (Makhija and Gatterbauer, 2023) that this problem remains polynomial-time solvable with multiplicities (see Theorem 7.18, Proposition 8.20 and Example 8.22).
@@
Consider the computational problem of finding tuples to be removed from the input database ${\mathfrak{A}}$ so that ${\mathfrak{A}}\not\models\mu$. We observe that if the resilience problem (with or without multiplicities) for a union $\mu$ of conjunctive queries is in P, then this problem also is in P. To see this, let $u\in\mathbb{N}$ be threshold. If $u=0$, then no tuple needs to be found and we are done. Otherwise, for every tuple $t$ in a relation $R^{{\mathfrak{A}}}$, we remove all copies of $t$ from $R^{{\mathfrak{A}}}$ and test the resulting database with the threshold $u-m$, where $m$ is the multiplicity of $t$. If the modified instance is accepted, then $t$ is a correct tuple to be removed and we may proceed to find a solution of this modified instance, otherwise we go one step back and try to remove a different tuple.
@@ 8.1. Connectivity
We show that when classifying the resilience problem for conjunctive queries, it suffices to consider queries that are connected. A $\tau$-structure is connected if it cannot be written as the disjoint union of two $\tau$-structures with non-empty domains.
@@
All terminology introduced for $\tau$-structures also applies to conjunctive queries with signature $\tau$: by definition, the query has the property if the canonical database has the property.
@@
We first present a polynomial-time reduction from the resilience problem of $\nu_{i}$, for some $i\in\{1,\dots,k\}$, to the resilience problem of $\nu$. Given an instance ${\mathfrak{A}}$ of the resilience problem for $\nu_{i}$, let $m$ be the number of tuples in relations of ${\mathfrak{A}}$. Let ${\mathfrak{A}}^{\prime}$ be the disjoint union of ${\mathfrak{A}}$ with $m$ copies of the canonical database of $\nu_{j}$ for every $j\in\{1,\dots,k\}\setminus\{i\}$. Observe that ${\mathfrak{A}}^{\prime}$ can be computed in polynomial time in the size of ${\mathfrak{A}}$ and that the resilience of ${\mathfrak{A}}$ with respect to $\nu_{i}$ equals the resilience of ${\mathfrak{A}}^{\prime}$ with respect to $\nu$.
@@ 8.2. Finite Duals
If $\mu$ is a union of conjunctive queries with signature $\tau$, then a dual of $\mu$ is a $\tau$-structure ${\mathfrak{B}}$ with the property that a finite $\tau$-structure ${\mathfrak{A}}$ has a homomorphism to ${\mathfrak{B}}$ if and only if ${\mathfrak{A}}$ does not satisfy $\mu$. The conjunctive query in Example 8.2, for instance, even has a finite dual. There is an elegant characterisation of the (unions of) conjunctive queries that have a finite dual. To state it, we need some basic terminology from database theory.
@@
The incidence graph of a relational $\tau$-structure ${\mathfrak{A}}$ is the bipartite undirected multigraph whose first colour class is $A$, and whose second colour class consists of expressions of the form $R(b)$ where $R\in\tau$ has arity $k$, $b\in A^{k}$, and ${\mathfrak{A}}\models R(b)$. An edge $e_{a,i,R(b)}$ joins $a\in A$ with $R(b)$ if $b_{i}=a$. A structure is called incidence-acyclic (also known as Berge-acyclic) if its incidence graph is acyclic, i.e., it contains no cycles (if two vertices are linked by two different edges, then they establish a cycle). A structure is called a tree if it is incidence-acyclic and connected in the sense defined in Section 8.1.
What follows is due to Nešetřil and Tardif (Nešetřil and Tardif, 2000); see also (Foniok, 2007; Larose et al., 2007).
@@
The theorem shows that in particular the query $\mu$ from Example 8.3 does not have a finite dual, since the query given there is not incidence-acyclic and hence cannot be homomorphically equivalent to a tree. To construct valued structures from duals, we introduce the following notation.
@@
Note that $\operatorname{Aut}({\mathfrak{B}})=\operatorname{Aut}(\Gamma({\mathfrak{B}},\sigma))$ for any $\tau$-structure ${\mathfrak{B}}$ and any $\sigma$. In the following result we use a correspondence between resilience problems for incidence-acyclic conjunctive queries and valued CSPs. The result then follows from the P versus NP-complete dichotomy theorem for valued CSPs over finite domains stated in Theorem 7.15.
@@
Theorem 8.11 can be combined with the tractability results for VCSPs from Section 7 that use fractional polymorphisms. To illustrate fractional polymorphisms and how to find them, we revisit a known tractable resilience problem from (Meliou et al., 2010; Freire et al., 2015a, b, 2020) and show that the corresponding valued structure has a fractional canonical pseudo cyclic polymorphism.
@@
There is a finite dual ${\mathfrak{B}}$ of $\mu$ with domain $\{0,1\}$ which is finitely bounded homogeneous, as described in Example 8.2. That example also describes a valued structure $\Gamma$ which is actually $\Gamma({\mathfrak{B}},\emptyset)$. Let $\omega$ be the fractional cyclic operation given by $\omega(\min)=\omega(\max)=\frac{1}{2}$. The fractional operation $\omega$ improves both valued relations $R$ and $S$ (they are submodular; see, e.g., (Krokhin and Živný, 2017)) and hence is a canonical cyclic fractional polymorphism of $\Gamma$. Combining Theorem 7.15 and 8.11, Example 8.12 reproves the results from (Freire et al., 2015b) (without multiplicities) and (Makhija and Gatterbauer, 2023) (with multiplicities) that the resilience problem for this query is in P.
In the following example we generalize the approach from Example 8.12.
@@
Let $\tau=\{R_{1},\dots,R_{n}\}$ be the signature of $\mu$. We describe a finite dual ${\mathfrak{B}}$ of $\mu$. The domain $B$ consists of $2^{n-1}$ elements $v_{S}$, which we index by a subset $S$ of $[n-1]$. Then the dual satisfies $R_{i}(v_{S},v_{T})$ for $S,T\subseteq[n-1]$ and $i\in\{1,\dots,n\}$ if and only if
@@
To see that ${\mathfrak{B}}$ is indeed a dual of $\mu$, let ${\mathfrak{A}}$ be a finite relational $\tau$-structure that does not satisfy $\mu$, and let $a\in A$. Let $S_{a}$ be the set of all elements $i\in[n-1]$ such that there exist elements $a_{i+1},\dots,a_{n}\in A$ such that
@@
Suppose that $a,b\in A$ are such that $R_{j}(h(a),h(b))$ does not hold in ${\mathfrak{B}}$ for some $j\in[n]$. This implies that
@@
First suppose that $1<i<n$. If the left-hand side in (13) is equal to $2$, then the disequality holds trivially. Whenever the left-hand side in (13) is equal to $1$, we must have $(i-1)\in S$, $(i-1)\in P$, $i\notin T$, or $i\notin Q$. Hence $(i-1)\in S\cup P$ or $i\notin T\cap Q$. In the first case, we have $R_{i}(v_{S\cup P},v_{T\cup Q})=0$, and in the second case we have $R_{i}(v_{S\cap P},v_{T\cap Q})=0$, and in both cases (13) holds. Finally, if the left-hand side in (13) is 00, then $R_{i}(v_{S},v_{T})=R_{i}(v_{P},v_{Q})=0$. Thus we have $(i-1)\in S$ or $i\notin T$ and at the same time $(i-1)\in P$ or $i\notin Q$. Therefore, we have $(i-1)\in S\cap P$, $i\notin T\cup Q$, or both $(i-1)\in S\cup P$ and $i\notin T\cap Q$. It follows that $R_{i}(v_{S\cup P},v_{T\cup Q})=R_{i}(v_{S\cap P},v_{T\cap Q})=0$. The case that $i=1$ and the case that $i=n$ can be treated similarly. Therefore, $\omega$ is a cyclic fractional polymorphism of $\Gamma({\mathfrak{B}},\emptyset)$, which by Theorem 8.11 implies that the resilience problem for $\mu$ is in P.

### v6

@@ 8. Application: Resilience
Let $\tau$ be a finite relational signature and $\mu$ a conjunctive query over $\tau$. The input to the resilience problem for $\mu$ consists of a finite $\tau$-structure $\mathfrak{A}$}, called a ($\tau$-)database [^1] To be precise, a finite relational structure is not exactly the same as a database, because the latter may not contain elements that are not contained in any relation. This difference, however, is inessential for the problems studied in this article., and the task is to compute the number of tuples that have to be removed from relations of $\mathfrak{A}$} so that $\mathfrak{A}$} does not satisfy $\mu$. This number is called the resilience of $\mathfrak{A}$} (with respect to $\mu$). As usual, this can be turned into a decision problem where the input also contains a natural number $u\in{\mathbb{N}}$ and the question is whether the resilience is at most $u$. Clearly, $\mathfrak{A}$} does not satisfy $\mu$ if and only if its resilience is 0.
@@
A natural variation of the problem is that the input database is a bag database, meaning that it may contain tuples with multiplicities. Formally, a multiset relation on a set $A$ of arity $k$ is a multiset with elements from $A^{k}$ and a bag database $\mathfrak{A}$} over a relational signature $\tau$ consists of a finite domain $A$ and for every $R\in\tau$ of arity $k$, a multiset relation $R^{\mathfrak{A}}$ of arity $k$. A bag database $\mathfrak{A}$} satisfies a union of conjunctive queries $\mu$ if the relational structure obtained from $\mathfrak{A}$} by forgetting the multiplicities of tuples in its relations satisfies $\mu$. In this article, we focus on bag databases whose relevance has already been discussed in the introduction. The reader might wonder why, unlike in (Makhija and Gatterbauer, 2023), we restrict to study of resilience problems in bag semantics and do not consider set semantics. This will become apparent from the proof of Proposition 8.20, see Remark 8.21.
@@
The basic resilience problem defined above can be generalized by admitting the decoration of databases with a subsignature $\sigma\subseteq\tau$, in this way declaring all tuples in $R^{{\mathfrak{A}}}$, $R\in\sigma$, to be exogenous. This means that we are not allowed to remove such tuples from $\mathfrak{A}$} to make $\mu$ false; the tuples in the other relations are then called endogenous. For brevity, we also refer to the relations in $\sigma$ as being exogenous and those in $\tau\setminus\sigma$ as being endogenous. If not specified, then $\sigma=\emptyset$, i.e., all tuples are endogenous. As an alternative, one may also declare individual tuples as being endogenous or exogenous. Under bag semantics, however, this case can be reduced to the one studied here (see Remark 8.22). The resilience problem that we study is summarized in Figure 1. Note that for every $\mu$ this problem is in NP. We remark that in Section 8.3, we briefly deviate from our focus on unions of conjunctive queries and, by adapting our results from Section 8.2, we establish a complexity dichotomy for resilience problems for two-way regular path queries; all the necessary definitions will be given in Section 8.3.
Fixed: a relational signature $\tau$, a subset $\sigma\subseteq\tau$, and a union $\mu$ of conjuctive queries over $\tau$.
Input: A bag database $\mathfrak{A}$} in signature $\tau$ and $u\in\mathbb{N}$.
$m:=$ minimal number of tuples to be removed from the relations in $\{R^{{\mathfrak{A}}}\mid R\in\tau\setminus\sigma\}$ so that ${\mathfrak{A}}\not\models\mu$.
Output: Is $m\leq u$?
Figure 1. The resilience problem considered in this article.:
@@
Observe that a finite $\tau$-structure satisfies $\mu$ if and only if it does not have a homomorphism to the $\tau$-structure $\mathfrak{B}$} with domain $B=\{0,1\}$ and the relations $R^{{\mathfrak{B}}}=\{(0,1),(1,1)\}$ and $S^{{\mathfrak{B}}}=\{(0,0),(0,1)\}$ (see Figure 2). We turn $\mathfrak{B}$} into the valued structure $\Gamma$ with domain $\{0,1\}$ where $R^{\Gamma}(0,1)=R^{\Gamma}(1,1)=0=S^{\Gamma}(0,0)=S^{\Gamma}(0,1)$ and $R^{\Gamma}$ and $S^{\Gamma}$ take value 1 otherwise. Then $\operatorname{VCSP}(\Gamma)$ is precisely the resilience problem for $\mu$ (with multiplicities). In Example 8.12 we reprove the result from (Makhija and Gatterbauer, 2023) that even with multiplicities, the problem can be solved in polynomial time.
Figure 2. The query $\mu$ from Example 8.2 (on the left) and the corresponding structure $\mathfrak{B}$} (on the right).:
@@
This query is linear in the sense of Freire, Gatterbauer, Immerman, and Meliou and thus its resilience problem without multiplicities can be solved in polynomial time (Theorem 4.5 in (Meliou et al., 2010); also see Fact 3.18 in (Freire et al., 2015a)). In Example 8.24 we reprove the result from (Makhija and Gatterbauer, 2023) that this problem remains polynomial-time solvable with multiplicities.
@@
Consider the computational problem of finding a set of tuples to be removed from the input database $\mathfrak{A}$} so that ${\mathfrak{A}}\not\models\mu$. We observe that if the resilience problem (with or without multiplicities) for a union $\mu$ of conjunctive queries is in P, then this problem also is in P. To see this, let $u\in\mathbb{N}$ be threshold. If $u=0$, then no tuple needs to be found and we are done. Otherwise, for every tuple $t$ in a relation $R^{{\mathfrak{A}}}$, we remove all copies of $t$ from $R^{{\mathfrak{A}}}$ and test the resulting database with the threshold umu-m, where $m$ is the multiplicity of $t$. If the modified instance is accepted, then $t$ is a correct tuple to be removed and we may proceed to find a solution of this modified instance, otherwise we go one step back and try to remove a different tuple.
@@ 8.1. Connectivity
We show that when classifying the resilience problem for conjunctive queries, it suffices to consider queries that are connected. This was observed already in (Freire et al., 2015b) for conjunctive queries in the set semantics setting; we provide a proof in our setting for completeness. A $\tau$-structure is connected if it cannot be written as the disjoint union of two $\tau$-structures with non-empty domains.
@@
All terminology introduced for $\tau$-structures also applies to conjunctive queries (i.e., primitive positive formulas) with signature $\tau$: by definition, the query has the property if the canonical database has the property.
@@
We first present a polynomial-time reduction from the resilience problem of $\nu_{i}$, for some $i\in\{1,\dots,k\}$, to the resilience problem of $\nu$. Given an instance $\mathfrak{A}$} of the resilience problem for $\nu_{i}$, let $m$ be the total multiplicity of tuples in relations of $\mathfrak{A}$}. Let $\mathfrak{A}$$^{\prime}$ be the disjoint union of $\mathfrak{A}$} with $m$ copies of the canonical database of $\nu_{j}$ for every $j\in\{1,\dots,k\}\setminus\{i\}$. Observe that $\mathfrak{A}$$^{\prime}$ can be computed in polynomial time in the size of $\mathfrak{A}$} and that the resilience of $\mathfrak{A}$} with respect to $\nu_{i}$ equals the resilience of $\mathfrak{A}$$^{\prime}$ with respect to $\nu$.
@@ 8.2. Finite Duals
If $\mu$ is a union of conjunctive queries with signature $\tau$, then a dual of $\mu$ is a $\tau$-structure $\mathfrak{B}$} with the property that a finite $\tau$-structure $\mathfrak{A}$} has a homomorphism to $\mathfrak{B}$} if and only if $\mathfrak{A}$} does not satisfy $\mu$. The conjunctive query in Example 8.2, for instance, even has a finite dual. There is an elegant characterization of the (unions of) conjunctive queries that have a finite dual. To state it, we need some basic terminology from database theory.
@@
The incidence graph of a relational $\tau$-structure $\mathfrak{A}$} is the bipartite undirected multigraph whose first color class is $A$, and whose second color class consists of expressions of the form $R(b)$ where $R\in\tau$ has arity $k$, $b\in A^{k}$, and ${\mathfrak{A}}\models R(b)$. An edge $e_{a,i,R(b)}$ joins $a\in A$ with $R(b)$ if $b_{i}=a$. A structure is called incidence-acyclic (also known as Berge-acyclic) if its incidence graph is acyclic, i.e., it contains no cycles (if two vertices are linked by two different edges, then they establish a cycle). A structure is called a tree if it is incidence-acyclic and connected in the sense defined in Section 8.1.
Note that every relational $\tau$-structure that is homomorphically equivalent to a tree is incidence-acyclic, but the reverse implication does not hold. However, the property of being homomorphically equivalent to a tree is weaker than being a tree. What follows is due to Nešetřil and Tardif (Nešetřil and Tardif, 2000); see also (Foniok, 2007; Larose et al., 2007).
@@
The theorem shows that in particular the query $\mu$ from Example 8.3 does not have a finite dual, since the query given there is not incidence-acyclic. To construct valued structures from duals, we introduce the following notation.
@@
Note that $\operatorname{Aut}({\mathfrak{B}})=\operatorname{Aut}(\Gamma({\mathfrak{B}},\sigma))$ for any $\tau$-structure $\mathfrak{B}$} and any $\sigma$. In the following result we use the correspondence between resilience problems for incidence-acyclic conjunctive queries and valued CSPs introduced in the present article. The result then follows from the P versus NP-complete dichotomy theorem for valued CSPs over finite domains stated in Theorem 7.15.
@@
Theorem 8.11 can be combined with the tractability results for VCSPs from Section 7 that use fractional polymorphisms to obtain tractability results for concrete resilience problems. To illustrate fractional polymorphisms and how to find them, we revisit a known tractable resilience problem from (Meliou et al., 2010; Freire et al., 2015a, b, 2020) and show that the corresponding valued structure has a fractional canonical pseudo cyclic polymorphism.
@@
There is a finite dual $\mathfrak{B}$} of $\mu$ with domain $\{0,1\}$ which is finitely bounded homogeneous, as described in Example 8.2. That example also describes a valued structure $\Gamma$ which is actually $\Gamma({\mathfrak{B}},\emptyset)$. Let $\omega$ be the fractional cyclic operation given by $\omega(\min)=\omega(\max)=\frac{1}{2}$. The fractional operation $\omega$ improves both valued relations $R$ and $S$ (they are submodular; see, e.g., (Krokhin and Živný, 2017)) and hence is a cyclic fractional polymorphism of $\Gamma$. Combining Theorems 7.15 and 8.11, Example 8.12 reproves the results from (Freire et al., 2015b) (without multiplicities) and (Makhija and Gatterbauer, 2023) (with multiplicities) that the resilience problem for this query is in P.
In the following example we generalize the approach from Example 8.12; tractability for this class of resilience problems in set semantics was proved in (Freire et al., 2015b) and in bag semantics in (Makhija and Gatterbauer, 2023).
@@
Let $\tau=\{R_{1},\dots,R_{n}\}$ be the signature of $\mu$. We describe a finite dual $\mathfrak{B}$} of $\mu$. The domain $B$ consists of $2^{n-1}$ elements $v_{S}$, which we index by a subset $S$ of $\{1,\dots,n-1\}$. Then the dual satisfies $R_{i}(v_{S},v_{T})$ for $S,T\subseteq\{1,\dots,n-1\}$ and $i\in\{1,\dots,n\}$ if and only if
@@
To see that $\mathfrak{B}$} is indeed a dual of $\mu$, let $\mathfrak{A}$} be a finite relational $\tau$-structure that does not satisfy $\mu$, and let $a\in A$. Let $S_{a}$ be the set of all elements $i\in\{1,\dots,n-1\}$ such that there exist elements $a_{i+1},\dots,a_{n}\in A$ such that
@@
Suppose that $a,b\in A$ are such that $R_{j}(h(a),h(b))$ does not hold in $\mathfrak{B}$} for some $j\in\{1,\dots,n\}$. This implies that
@@
First suppose that $1<i<n$. If the left-hand side in (13) is equal to $2$, then the disequality holds trivially. Whenever the left-hand side in (13) is equal to $1$, we must have $(i-1)\in S$, $(i-1)\in P$, $i\notin T$, or $i\notin Q$. Hence $(i-1)\in S\cup P$ or $i\notin T\cap Q$. In the first case, we have $R_{i}(v_{S\cup P},v_{T\cup Q})=0$, and in the second case we have $R_{i}(v_{S\cap P},v_{T\cap Q})=0$, and in both cases (13) holds. Finally, if the left-hand side in (13) is 0, then $R_{i}(v_{S},v_{T})=R_{i}(v_{P},v_{Q})=0$. Thus we have $(i-1)\in S$ or $i\notin T$ and at the same time $(i-1)\in P$ or $i\notin Q$. Therefore, we have $(i-1)\in S\cap P$, $i\notin T\cup Q$, or both $(i-1)\in S\cup P$ and $i\notin T\cap Q$. It follows that $R_{i}(v_{S\cup P},v_{T\cup Q})=R_{i}(v_{S\cap P},v_{T\cap Q})=0$. The case that $i=1$ and the case that $i=n$ can be treated similarly. Therefore, $\omega$ is a cyclic fractional polymorphism of $\Gamma({\mathfrak{B}},\emptyset)$, which by Theorem 8.11 and Theorem 7.15 implies that the resilience problem for $\mu$ is in P.